OpenStudy (anonymous):

Hey I have another one.. can you please help me on how to graph this? Thanks so much!! x2- 3x greater than or equal to 0

OpenStudy (anonymous):

so you want all the values of x for which the graph of y = x^2 - 3*x is above the x-axis or touching...so it helps to graph that parabola... if you picture the parabola in your head , you can see it will be for x is greater than or equal to 3 or x is smaller than or equal to 0

OpenStudy (anonymous):

would it be a solid line?

OpenStudy (anonymous):

when you graph the parabola

OpenStudy (anonymous):

nope...whenever you sketch a parabola, it helps to get three pieces of information: a) where it touches the x-axis, where it touches the y-axis, and where its turning point is. to find the x-intercepts, you make y =0: x^3 - 3x = 0, so x(x - 3) = 0, so x = 0 or x = 3 so the x-intercepts are (3,0) and (0,0) the y-intercept is (0,0) you should now have enough information to sketch it

OpenStudy (anonymous):

thank you i got it!